# LEREL: Lipschitz Continuity-Constrained Emotion Recognition Ensemble Learning For Electroencephalography

### v1

@@ 3.5 Validation on the EAV Passive Dataset
The EAV dataset includes both active (recorded during conversations) and passive (collected during stimulus listening) signals. Passive signals present greater analytical challenges due to their higher noise levels, attenuated amplitude and weaker correlations with affective states. To evaluate model robustness, we specifically tested performance on passive signals. As shown in Figure 10, LEREL maintained a high classification accuracy under these suboptimal conditions, demonstrating exceptional signal processing stability.
Figure 10: Using passive data experiments to compare with other active data experiments on the EAV dataset
Figure 11: Partitioning with contextual information
Figure 12: Partitioning without contextual information
Figure 13: Real-time validation test of EAV
Figure 14: Real-time validation test of FACED
@@ 3.6 Real-time Validation Test
The proposed LEREL model demonstrates robust real-time emotion recognition, achieving consistently high classification accuracy on both EAV and FACED datasets (Figures 14 and 14). These results confirm its stronggeneralization capacity and practical applicability in complex real-world environments. Specifically, LEREL’s temporal dependency modeling enables effective EEG feature extraction and noise suppression, ensuring stable performance under challenging conditions.
@@ 4 Conclusion
This paper presents LEREL, a robust EEG-based emotion recognition framework that integrates ensemble learning with Lipschitz regularization. The framework combines: 1) Lipschitz constraints for noise-robust feature extraction, and 2) ensemble classifiers to optimize bias-variance trade-offs, ensuring reliable generalization. Experimental results demonstrate LEREL’s superior performance in both offline and real-time scenarios, particularly with challenging EEG signals. Comprehensive evaluations across multiple datasets and ablation studies validate each component’s contribution, showing consistent high accuracy while maintaining stability under noisy conditions. These results highlight LEREL’s practical potential for real-world affective computing. Future work will investigate computational optimization and multimodal integration to further enhance emotion recognition capabilities.

### v2

@@ 4.6 Validation on the EAV Passive Dataset
The previous experimental results on the EAV positive dataset provide compelling evidence of the superiority and robustness of the LEL framework. These results demonstrate that LEL consistently delivers stable and reliable performance in emotion recognition tasks, even when applied to diverse signal types. The EAV dataset comprises both active and passive signals, with the latter being particularly challenging due to their high noise levels, low amplitudes, and weak correlations with emotional states.
To rigorously evaluate LEL’s robustness, we applied the model specifically to passive signals. As shown in Figure 7, LEL maintained high classification accuracy under these adverse conditions, highlighting its strong signal processing stability. Notably, the LEL variant with Lipschitz constraints (Lc – ECL Best) consistently achieved the highest accuracy across most emotion recognition tasks. Even in instances where other models outperformed LEL in specific cases, its performance remained comparably high, with minimal degradation.
These findings underscore LEL’s effectiveness not only in handling passive signals but also in delivering reliable performance across a broad range of emotion recognition scenarios.
Figure 7: Using passive data experiments to compare with other active data experiments on the EAV dataset.
@@ 4.7 Real-time Validation Test
The proposed LEL model demonstrates robust real-time emotion recognition, achieving consistently high classification accuracy on EAV, FACED and SEED datasets (Figures 10). These results confirm its stronggeneralization capacity and practical applicability in complex real-world environments. Specifically, LEL’s temporal dependency modeling enables effective EEG feature extraction and noise suppression, ensuring stable performance under challenging conditions.
Figure 8: Partitioning with contextual information.
Figure 9: Partitioning without contextual information.
Figure 10: Real-time validation test on EAV (a), FACED (b) and SEED (c)
4.8 Correlation Between EEG Channels and Emotions
An analysis of the nine emotions in the FACED dataset shows that the channels with high - impact proportions for different emotions vary significantly. To better illustrate the relationships between these channels, chord diagrams are used, where the connection strength between channels is represented by the depth of color in the chords, offering a visual depiction of their interrelations.
These chord diagrams reveal the unique neural patterns associated with different emotions, highlighting their distinct activity modes in the brain (Figure 11). The findings are highly significant for understanding the brain mechanisms of emotion processing and for developing EEG-based emotion recognition systems. By identifying channel combinations linked to specific emotions, feature extraction becomes more efficient, which in turn enhances the accuracy of emotion classification.
Figure 11: This figure shows nine chord diagrams representing different emotions: Amusement, Anger, Disgust, Fear, Inspiration, Joy, Neutral, Sadness and Tenderness. Each diagram illustrates the connections and relationships between various elements or features associated with the respective emotion.
5 Discussion
The experimental results underscore LEL’s effectiveness in offline and real-time emotion recognition, particularly in handling noisy EEG signals. This stems from two key elements: Lipschitz constraints for robust feature extraction and ensemble classifiers for optimizing bias-variance trade-offs. These components work together to reduce noise impact, accelerate convergence, enhance stability, and improve generalization by combining multiple classifier predictions.
LEL’s performance was validated through extensive experiments across the EAV, FACED, and SEED datasets. On the EAV positive dataset, it delivered an average accuracy of 76.43% and an F1-score of 76.22%, surpassing existing methods and showing balanced emotion category recognition. When contrasted with AMERL’s EEG benchmark, LEL exhibited significant advantages in multiple metrics, especially in EEG data processing. In the FACED dataset experiments, LEL achieved 83.00% average accuracy, outperforming baseline metrics and demonstrating its ability to handle heterogeneous EEG signals with high-precision emotion recognition. On the SEED dataset, it obtained 87.22% accuracy and 87.06% F1-score, proving its capability in complex emotion recognition tasks and adaptability across different paradigms, alongside strong robustness to EEG signal noise and interference. These results collectively demonstrate LEL’s effectiveness and robustness in EEG-based emotion recognition.
Through experiments analyzing the training process, ROC curves, and t-SNE visualizations, it was found that Lipschitz constraints significantly enhance training stability, especially on low-accuracy, noisy datasets with complex feature distributions. The model’s emotion classification ability improves with increasing training epochs, and the boundaries between different emotions become clearer with appropriate Lipschitz constants. Ablation studies further confirmed the effectiveness of the ensemble strategy, as each individual classifier achieved high accuracy, and the integrated model showed significant improvements in average accuracy and F1-score.
LEL’s robustness was also validated on the EAV passive dataset, where it maintained high classification accuracy despite high noise, low amplitude, and weak emotional correlations, confirming its excellent signal processing stability. Real-time validation tests further demonstrated LEL’s strong generalization and practical applicability, achieving consistently high classification accuracy on both the EAV and FACED datasets.
Moreover, experiments analyzing channel relationships using chord diagrams revealed unique neural patterns associated with different emotions, providing valuable insights into the brain mechanisms of emotion processing and offering a foundation for the development of EEG-based emotion recognition systems. These findings collectively support LEL’s advantages and potential in the field of EEG-based emotion recognition.
@@ 6 Conclusion
In this paper, we introduced LEL (Lipschitz continuity-constrained Ensemble Learning), a novel framework that significantly advances the field of EEG-based emotion recognition. By integrating Lipschitz regularization with ensemble learning strategies, LEL effectively addresses key limitations in existing EER methods, including model instability, poor generalization, and vulnerability to noise and intra-subject variability. LEL demonstrates superior performance across multiple benchmark datasets, underscoring the framework’s robustness and reliability, particularly in challenging scenarios involving noisy or low-amplitude EEG signals. The application of Lipschitz constraints enhances model stability and generalization, while ensemble learning reduces bias and variance by leveraging multi-classifier fusion. Beyond its empirical success, LEL holds strong potential for real-world affective computing applications, where accurate and efficient emotion recognition is essential for enabling intuitive human-computer interactions. Future work will focus on optimizing computational efficiency and exploring multimodal data integration to further improve recognition performance and expand the framework’s applicability.
CRediT authorship contribution statement
Shengyu Gong: Writing - original draft, Methodology, Formal analysis, Conceptualization, Visualization, Software; Yueyang Li: Writing - original draft, Investigation, Methodology; Zijian Kang: Validation, Data Curation, Investigation; Weiming Zeng: Resources, Supervision, Writing - Review & Editing; Hongjie Yan: Validation, Writing - Review & Editing; Zhiguo Zhang: Validation, Writing - Review & Editing; Wai Ting Siok: Funding acquisition, Validation, Writing - Review & Editing; Nizhuan Wang: Conceptualization, Writing - review & editing, Supervision, Project administration, Funding acquisition.
Acknowledgments
This work was supported by The Hong Kong Polytechnic University Start-up Fund (Project ID: P0053210), The Hong Kong Polytechnic University Faculty Reserve Fund (Project ID: P0053738), an internal grant from The Hong Kong Polytechnic University (Project ID: P0048377), The Hong Kong Polytechnic University Departmental Collaborative Research Fund (Project ID: P0056428), The Hong Kong Polytechnic University Collaborative Research with World-leading Research Groups Fund (Project ID: P0058097) and Research Grants Council Collaborative Research Fund (Project ID: P0049774).